40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 12 08:40:06 2016, Steve, Update, PEM, Air cond maintenance 
    Reply  Mon Jan 25 11:06:08 2016, Steve, Update, PEM, noisy fan belt replaced 
    Reply  Fri Jan 29 08:18:13 2016, Steve, Update, PEM, PEM PEM.png
       Reply  Wed Apr 20 14:36:45 2016, Steve, Update, PEM, dusty lab AP_accm.jpgwiped_AP.jpgtemp-partic_6months.png3_years_PEM.png
          Reply  Tue Jun 21 14:12:10 2016, Steve, Update, PEM, vacuum envelope wiped 60dPEM.png
       Reply  Fri Jan 13 14:18:28 2017, Steve, Update, PEM, air condition fixed PEM_120d.png
Message ID: 12712     Entry time: Fri Jan 13 14:18:28 2017     In reply to: 11959
Author: Steve 
Type: Update 
Category: PEM 
Subject: air condition fixed 

The old control room AC  has been stick in heating mode for about 2 months. It's thermostate and fan belt  was finally replaced. It was calibrated and set to 71 F ( just behind 1X6 on west wall ) around 1pm.

Out belt; sad inside 

at 4 pm Rana cried

It must be too tight.

Attachment 1: PEM_120d.png  46 kB  Uploaded Wed Jan 18 08:45:07 2017  | Hide | Hide all
PEM_120d.png
ELOG V3.1.3-